OpenStudy (anonymous):

Medal!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!Write an equation of a line parallel to y = 1 over 4x − 6 in slope-intercept form that passes through the point (−1, 5). y= 1 over 4x − 1 y= 1 over 4x + 5.25 y= −4x + 5 y= −4x − 1.25

OpenStudy (anonymous):

anyone good at this

OpenStudy (here_to_help15):

Let me see hmmm.

OpenStudy (anonymous):

i suck at slope intercept forms on a grid

OpenStudy (anonymous):

@Here_to_Help15

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@iambatman

OpenStudy (igreen):

Parallel to this? \(y = \dfrac{1}{4x-6}\)

OpenStudy (anonymous):

yes

OpenStudy (phi):

I am sure they gave you the equation of a line in slope-intercept form: y = ¼ x − 6 and want a line parallel to this one, but goes through the point (-1,5)

OpenStudy (anonymous):

this is hard for me

OpenStudy (anonymous):

yes

OpenStudy (phi):

The first thing to notice is the line they gave you is in slope-intercept form y = mx + b where m is the slope. what is the slope of the line they gave you ? what is the slope of y = ¼ x − 6 ?

OpenStudy (anonymous):

@phi how do you get slope from that

OpenStudy (anonymous):

@NinjaDevo

OpenStudy (phi):

you match up the equation with y = m x + b m is the slope the number in front of the x is the slope

OpenStudy (anonymous):

so its y=1/4x+6

OpenStudy (phi):

all we want (so far) is the slope. what is the slope of y = ¼ x − 6 ?

OpenStudy (phi):

the number in front of the x is the slope

OpenStudy (anonymous):

how do you get it?

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

all i know is how to write it i dont know how to solve that

OpenStudy (anonymous):

1/4 is the slope

OpenStudy (phi):

yes. ¼ is the slope. next, we want an equation that looks like y = m x + b where m is the slope. we know the slope is ¼ (the same slope as the line they gave us) so replace m with ¼ y = ¼ x + b now we need to find b. To find b, we use the info passes through the point (−1, 5). that means when x = -1 y is 5 use those numbers in y = ¼ x + b can you replace y with 5 and x with -1 ?

OpenStudy (anonymous):

5=1/4-1+b

OpenStudy (phi):

ok. except ¼ x is short-hand for ¼ times x when you put in -1, it means ¼ * -1 so you should write it 5=¼ *-1+b -1*¼ is -¼ so we can write it 5 = -¼ + b now add +¼ to both sides

OpenStudy (anonymous):

i dont understand that?

OpenStudy (anonymous):

can you write that part out for me

OpenStudy (anonymous):

im sorry im just really bad at geometry

OpenStudy (phi):

luckily this is algebra. do you get 5=¼ *-1+b

OpenStudy (anonymous):

yes

OpenStudy (phi):

-1 * ¼ is - ¼ (just like -1*2= -2 and -1 * ½ is -½ , and so on)

OpenStudy (anonymous):

ok i understand

OpenStudy (phi):

we have 5=-¼ +b

OpenStudy (phi):

add +¼ to both sides 5+¼ =-¼ + ¼ + b on the right side -¼ + ¼ is ?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

so its 5+1/4=b

OpenStudy (phi):

so we have 5 + ¼ = 0+b 0 + anything is anything (adding 0 does not change it). so 0+b is just b 5+¼ = b we can write 5+¼ as a mixed number 5 ¼ (means the same thing) or change ¼ to a decimal 0.25 and write 5+0.25 = 5.25 we find 5.25 = b that means b is 5.25 use that in our equation (remember it ? ) y = ¼ x + b we get the answer y = ¼ x + 5.25

OpenStudy (anonymous):

ok thank you so much

OpenStudy (anonymous):

@phi can you help me with another?

OpenStudy (anonymous):

Find the perimeter of the shape below: A four sided figure RSTU is shown. R is located at negative 1, 3. S is located at negative 2, 7. T is located at 2, 5. U is located at negative 1, 5 7.4 units 8.9 units 11.2 units 13.6 units

OpenStudy (phi):

it's easier to read if you use - instead of negative.

OpenStudy (anonymous):

ok my bad

OpenStudy (phi):

it looks like you need to use the distance formula four times.

OpenStudy (anonymous):

ok

OpenStudy (phi):

does the picture look like this?

OpenStudy (anonymous):

yes

OpenStudy (phi):

the distance formula is \[d= \sqrt{ (x_1-x_2)^2 + (y_1-y_2)^2 } \] between points \( (x_1,y_1) \text{ and } (x_2,y_2) \)

OpenStudy (phi):

R is (-1,3) S is (-2,7) can you put those numbers into the formula?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

its comes out to be the sqrt of 17

OpenStudy (anonymous):

i plugged it in and solved and thats what i got

OpenStudy (phi):

ok. But all the choices are decimal, so we should write it as (about) 4.12 (we keep 2 decimals and will round to 1 decimal at the end) one down, 3 to go now the distance between S is (-2,7) T is ( 2,5)

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

sqrt of 20

OpenStudy (phi):

as a decimal?

OpenStudy (anonymous):

4.47

OpenStudy (anonymous):

when its rounded

OpenStudy (phi):

now the last 2 sides T is ( 2,5) U is (-1, 5) (notice we don't need the distance formula because the y does not change. just count the x difference)

OpenStudy (anonymous):

ok so then its sqrt 9

OpenStudy (anonymous):

which is 3

OpenStudy (phi):

notice the difference is the x values -1 to 2 is 3 (simpler!)

OpenStudy (phi):

finally U is (-1, 5) R is (-1,3)

OpenStudy (anonymous):

@phi i only have like maybe 40 mins to turn this in can you give answers and then explanations just so i can get this done i know you want me to engage and ill try but i cant spend this amount of time on one question

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

ok its sqrt of 4 which is 2

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@myininaya

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@thomaster

OpenStudy (phi):

yes. now add up the 4 numbers ... the perimeter is the distance around the figure i.e. the sum of its sides

OpenStudy (phi):

and round to the nearest tenth

OpenStudy (anonymous):

ok its 13.6

OpenStudy (anonymous):

can you help me with some more please could yoiu give me the answer and then explain that way i can get this done

OpenStudy (anonymous):

i dont have much time and all i need is an explanation after and ill get it